OpenStudy (anonymous):

solve the equation 12-5x+9x=-2(-7x+9)

OpenStudy (lgbasallote):

distriibute two into your quantity then combine like terms :D dont forget to put all the x on one side and the constant on the other

OpenStudy (anonymous):

i got x= 1.6 but it's wrong because when i go to plug in the numbers it's not right-

OpenStudy (anonymous):

i don't know where i'm messing up at

OpenStudy (anonymous):

x=3

OpenStudy (lgbasallote):

12 - 5x + 9x = 14x - 18 did you get that same equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

12-4x=14x-18

OpenStudy (lgbasallote):

12 + 18 = 14x + 4x 30 = 18x

OpenStudy (lgbasallote):

yup 1.6 is right

OpenStudy (anonymous):

12-5x+9x=-2(-7x+9) 12-5x+9x=14x-18 12+4x=14x-18 -12 -12 4x=14x-30 -14x-14x -10x=-30 -x=-3 x=3

OpenStudy (anonymous):

thanks- i c where i messing up at now

OpenStudy (anonymous):

Check: 12-5(3)+9(3)=-2(-7(3)+9) 12-15+27=-2(-21+9) 12-15+27=42-18 -3+27=42-18 24=24

OpenStudy (anonymous):

First of all ,in such questions we must see whether it is an equation/inequation . Yours is a linear equation Linear equations are solved as: First bring the like terms together i.e. huge\12-5x+9x=-2(-7x+9) can be written as (with proper sign changes) 12-15+27=-2(-21+9) 12-15+27=42-18 -3+27=42-18 24=24